# Communication-Efficient Soft Actor-Critic Policy Collaboration via Regulated Segment Mixture in Internet of Vehicles

### v2

@@ V-A1 Segment Request & Response
Inspired by segmented pulling synchronization mechanism in DFL [69], we develop and perform a segment request & response procedure. The proposed approach divides transmission of the policy parameters into segments, and each agent selectively request various segments of policy parameters from different neighbors simultaneously through V2V communication, thereby facilitating the construction of a reconstructed referential policy for subsequent aggregation while also effectively balancing the load of communication costs and optimizing bandwidth usage. Specifically, for every communication round, each agent $i$ breaks its policy parameters $\theta^{(i)}$ into $P$ ($P\leq|\Omega_{i}|$, according to the number of different agents’ neighbors in current time) non-overlapping segments $\hat{\theta}^{(i)}_{1},\hat{\theta}^{(i)}_{2},\cdots,\hat{\theta}^{(i)}_{P}$ as
$\theta^{(i)}=(\hat{\theta}^{(i)}_{1},\hat{\theta}^{(i)}_{2},\cdots,\hat{\theta}^{(i)}_{P})\vspace{-0.1em}$ (16)
@@
Significantly, the available segmentation strategies are diverse and include, but are not limited to, dividing the policy parameters by splitting DNN layers [38], modular approach [70], or segmenting according to the parameter size [69]. Meanwhile, there exist rather diverse V2V communication protocols and technologies such as DSRC (based on IEEE 802.11p and its subsequent IEEE 802.11bd), NR-V2X (as specified in 3GPP Release 16 and 17), which support Multiple Input Multiple Output (MIMO) technology and advanced signal processing techniques, such as spatial multiplexing and beamforming. Therefore, the communication capability depends on multiple factors ranging from physical layer configurations, such as antenna array setup and channel conditions, to upper layer protocols, shaping the theoretical upper limit for segmentation granularity $P_{\max}$. For instance, a $4\times 4$ MIMO system in a vehicle can theoretically support up to $4$ independent data streams of policy segments simultaneously, enhancing data throughput and robustness. Besides, shorter distances between vehicles and better channel conditions generally enhance the capacity and reliability of system. Still, the segmentation also can be dynamic, with larger $P_{\max}$ in environments with higher vehicle density and smaller $P_{\max}$ in cases where fewer vehicles are within communication range.
@@
Here, to clarify this process, we use the most intuitive uniform parameter partition. For each segment $p=1,\cdots,P$, agent $i$ randomly selects a target agent (without replacement) from its neighbors (i.e., $j_{p}\in\Omega_{i}$) to send segment request $(i,P,j_{p},p)$, which indicates the agent $i$ who initiates the request and its total segment number $P$, as well as the requested segment $p$ from the target agent $j_{p}$. Upon receiving the request, the agent $j_{p}$ will break its own policy parameters $\theta^{(j_{p})}$ into $P$ segments and return the corresponding requested segment $\hat{\theta}^{(j_{p})}_{p}$ according to the identifier $p$. The number of segments $P$ can be flexibly determined according to the actual communication range of agent $i$, i.e., $P=\min\{P,|\Omega_{i}|\}<P_{\max}$. It should be noted that in order to reduce the complexity and facilitate the implementation, we only discuss the case as in Fig. 2 that $P$ is the same constant for all agents and is not greater than $\max_{i}|\Omega_{i}|,\forall i$. Meanwhile, the segmentation methodology is identical across agents. Then, agent $i$ could reconstruct a referential policy based on all of the fetched segments, that is,
$\tilde{\theta}^{(i)}=(\hat{\theta}^{(j_{1})}_{1},\hat{\theta}_{2}^{(j_{2})},\cdots,\hat{\theta}_{P}^{(j_{P})})\vspace{-0.1cm}$ (17)
@@
In fact, this segmented transmission approach can be executed in parallel, thereby optimizing the utilization of available bandwidth. Instead of being confined to a single link, the traffic is distributed across $P$ links, enhancing the overall data transfer efficiency. Besides, in order to further accelerate the propagation and ensure the model quality, we can construct multiple model replicas in RSM-MASAC. That is, the process of segment request &\&& response can be repeated $P\times\kappa$ times, reconstructing $\kappa$ reconstructed referential policies in one communication round.
Algorithm 1 RSM-MASAC Algorithm
1:  Initialize network parameters $\theta^{(i)}$, $\omega_{1}^{(i)}$, $\omega_{2}^{(i)}$, $i=1,2,\cdots N$
2:  Initialize target network parameters $\bar{\omega}_{1}^{(i)}\leftarrow\omega_{1}^{(i)}$, $\bar{\omega}_{2}^{(i)}\leftarrow\omega_{2}^{(i)}$
@@
3:  Initialize learning rate $\eta_{\pi}$, $\eta_{Q}$, $\eta_{\alpha}$, temperature $\alpha$, communication interval $U$, number of segments $P$ and number of replicas $\kappa$
4:  Initialize iteration index $k\leftarrow 0$ and $counter\leftarrow 0$
@@
9:        Select an action $a_{t}^{(i)}$ with respect to $s_{t}^{(i)}$ according to the current policy $\pi^{(i)}$ parameterized by $\theta^{(i)}$
10:        Observe reward $r_{t}^{(i)}$ and next state $s_{t+1}^{(i)}$
11:        Save the new transition in replay buffer: $\mathcal{D}^{(i)}\leftarrow\mathcal{D}^{(i)}\cup\langle s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)}\rangle$
12:        Sample a mini-batch $\Phi^{(i)}\sim\mathcal{D}^{(i)}$
13:        Update soft $Q$ function $\omega_{x}^{(i)}\leftarrow\omega_{x}^{(i)}-\eta_{Q}\nabla J_{Q}(\omega_{x}^{(i)})$ by (6), for $\forall x={1,2}$
@@
15:           Update policy $\vspace{-0.5em}\theta_{k+1}^{(i)}\leftarrow\theta_{k}^{(i)}-\eta_{\pi}\nabla J_{\pi}(\theta_{k}^{(i)})$ by (7)
16:           Adjust temperature $\alpha^{(i)}\leftarrow\alpha^{(i)}-\eta_{\alpha}\nabla J(\alpha^{(i)})$ by (9)
17:           Update target networks $\bar{\omega}_{x}^{(i)}\leftarrow\varrho\omega_{x}^{(i)}+(1-\varrho)\bar{\omega}_{x}^{(i)}$, for $\forall x=1,2$
18:           $k\leftarrow k+1$
@@
20:        $counter\leftarrow counter+1$
@@
23:           Update policy $\theta_{k}^{(i)}\leftarrow$ CommMix($\theta_{k}^{(i)}$, $i$, $\Omega_{i}$, $P$, $\kappa$)
@@
27:  CommMix($\theta,i,\Omega_{i},P,\kappa$):
28:       for each replica $u=1,2,\cdots,\kappa$ do
29:           Send $P$ pulling request $(i,P,j_{p},p)$ to nearby collaborators   in $\Omega_{i}$, and receive $\hat{\theta}^{(j_{p})}_{p}$ to reconstruct $\tilde{\theta}$ as (17)
30:           Select $M$ samples from the replay buffer $\mathcal{D}^{(i)}$
31:           Estimate $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ according to (19)
32:           if $\mathbb{A}^{+}_{\pi}(\tilde{\pi})>0$ then
33:               Evaluate $F(\theta)$ according to (20)
34:               Get the upper bound of $\zeta$ according to Theorem 2
35:               Make the mixture metric $\zeta$ less than the calculated upper    bound, and update $\theta_{\text{mix}}^{(i)}$ by (10)
36:              $\quad\theta\leftarrow\theta_{\text{mix}}^{(i)}$
37:           end if
38:       end for
@@ V-A2 Policy Parameter Mixture with Theory-Established Performance Improvement
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\frac{1}{2}\zeta^{2}(\tilde{\theta}-\theta)^{\intercal}F(\theta)(\tilde{\theta}-\theta)\vspace{-2em}$
@@
$\vspace{-0.5em}F(\theta)=\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi}\\ a\sim\pi_{\theta}\end{subarray}}\left[\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)\!\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)^{\intercal}\right]$
@@
$\qquad+\alpha[H(\tilde{\pi}_{\tilde{\theta}}(\cdot|s_{t}))-H({\pi}_{{\theta}}(\cdot|s_{t}))]\bigg{]}$ (19)
@@
$F(\theta)\!\approx\!\mathop{\mathbb{E}}\limits_{s_{t}\sim\mathcal{D}}\!\bigg{[}\!\mathop{\mathbb{E}}\limits_{a_{t}\sim\pi_{\theta}}\!\bigg{[}\!\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)\!\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)^{\intercal}\!\bigg{]}\bigg{]}$ (20)
@@ V-B Communication Cost Analysis
Regarding the communication overhead for each segment request, RSM-MASAC incurs a data transmission cost of $\upsilon_{p}=\upsilon/P$ through V2V communications. Notably, we consider V2V communication as an implementation method without restricting it to specific protocols. Following the same settings as in [70], we provide illustrative examples to demonstrate the overall reduction in communication overhead. Specifically, the Collective Perception Message (CPM) defined in Collective Perception Service (CPS) of ETSI TR 103 562 [72] can encapsulate DNN parameters into the Perceived Object Containers (POCs) and propagate them among cooperating CAVs with each message carrying a payload of $4,480$ bytes, serving as segment responses in our algorithm. And the DNN parameters $\theta\in\mathbb{R}^{d}$ ’s transmission size $\upsilon$ can be regarded as $32d$ bits, commonly assumed in [62, 70, 24]. Consequently, the total communication overhead for reconstructing $\kappa$ referential policies per agent in each round $c(\upsilon,f)=\frac{N\times\upsilon\times\kappa}{8\times 1024^{3}}$ (GB), as well as $\frac{N\times\upsilon\times\kappa}{8\times 4480}$ message numbers, which is $(N-1)/\kappa$ times less than that in a fully connected communication setup [44]. Additionally, communication occurs at a periodic interval of $U$, allowing for further reduction in communication overhead by decreasing communication frequency. Moreover, by simultaneously requesting $P$ agents in parallel, RSM-MASAC benefits from the sufficient use of the bandwidth and enhances the capability to overcome possible channel degradation.
@@ V-C Complexity Analysis
As the calculation of policy advantage $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ depends on the sample size $M$ and parameter size $|d|$, with complexity $\mathcal{O}(M|d|)$. FIM calculation with sampling approximated method using first-order gradients in (20) scales with the number of parameters, with complexity $\mathcal{O}(M|d|^{2})$. Hence, with $\kappa$ replicas, the computation complexity of CommMix is $\mathcal{O}(\kappa M|d|^{2})$ and memory complexity is $\mathcal{O}(M|d|^{2})$. However, the actual computation of the FIM is not restricted to our sampling-based method alone. Considering models with a large number of parameters, there is extensive research on FIM approximation methods, such as Diagonal Approximation [73, 74] with $\mathcal{O}(\kappa M|d|r_{\text{rank}})$, Low-Rank Approximation [75] with $\mathcal{O}(\kappa M|d|)$, etc. Thus, the complexity can be further reduced.
@@ VI Experimental Results and Discussions
In this section, we validate the effectiveness of our proposed algorithm for CAVs speed control in IoV, highlighting its superiority compared to other methods.
@@ VI-A Experimental Settings
Parameters Definition Figure Eight Merge
@@
  • “Figure 8”: $14$ vehicles navigate a one-way lane shapes like a figure “8”, including $5$ emulated HDVs controlled by Simulation of Urban MObility (SUMO) with Intelligent Driver Model (IDM) [78], and $9$ IRL-controlled CAVs maintaining dedicated links to update their parameters through the V2V channel. At the lane’s intersection, each CAV adjusts its acceleration to traverse efficiently, aiming to boost the traffic flow’s average speed.
  • “Merge”: A highway on-ramp merging scenario with vehicle flow at $2,300$ per hour, including a maximum of $2,200$ vehicles on the main road and $100$ vehicles on the ramp. Within each epoch, $13$ vehicles are randomly chosen to instantiate the DRL-based controllers, aiming to manage collision avoidance and congestion at merge points. The simulation settings closely align with those of the first scenario.
Both two scenarios are modified to assign the limited partial observation of the global environment as the state of each CAV, including the position and speed of its own, the vehicle ahead and behind. Only CAVs can execute the V2V end-to-end communication, and the connectivity of V2V links at time $t$ depends on the CAVs’ position and communication range, which are both extracted using the TraCI simulator in our experiments. A communication range of $90$m is adopted in interactions “Figure 8” and $400$m in highway “Merge”. This is according to the conclusion in [79] that communication range at intersections will be extremely reduced compared with the conventional scenarios [80], and relevant service requirements[81]. Unless otherwise stated, our experiments are based on the common MARL V2V lossless ideal communication premise [30, 31, 32, 33]. Meanwhile, each CAV’s action is a continuous variable representing the speed acceleration or deceleration and is normalized between $[-1,1]$.
@@
Number of samples to evaluate $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ and $F(\theta)$ $M$ $50$
Learning rate of actor network $\eta_{\pi}$ $10^{-4}$
Learning rate of critic network $\eta_{Q}$ $3\times 10^{-3}$
@@
Number of segments $P$ 4
Number of model replicas $\kappa$ 3
@@
In order to reduce the occurrence of collisions and promote the traffic flow to the maximum desired speed, we take the normalized average speed of all vehicles at each timestep as the individual reward in each scenario [^2] Notably, we assume complete knowledge of individual vehicle speeds at each vehicle here. Beyond the scope of this paper, some value-decomposition methods like [82] can be further leveraged to derive a decomposed reward, so as to loosen such a strict requirement., which is assigned to each training vehicle after its action is performed. In addition, the current epoch will be terminated once a collision occurs or the max length of step $T$ in an epoch is reached.
We perform tests every $10$ epochs and calculate the average reward based on the accumulated rewards from $5$ independent testing epochs. Besides, all results are produced by taking the average of $5$ independent simulations. Moreover, as for the baseline, we take the vehicles controlled by the Flow IDM [78], which belongs to a typical car-following model incorporating extensive prior knowledge and indicates the pinnacle of performance achievable by the best centralized federated MARL algorithms [27]. Furthermore, the principal hyper-parameters used in simulations are listed in Table IV.

### v3

@@ V-A1 Segment Request & Response
Inspired by segmented pulling synchronization mechanism in DFL [66], we develop and perform a segment request & response procedure. The proposed approach divides transmission of the policy parameters into segments, and each agent selectively requests various segments of policy parameters from different neighbors simultaneously through D2D communication, thereby facilitating the construction of a reconstructed referential policy for subsequent aggregation while also effectively balancing the load of communication costs and optimizing bandwidth usage. Specifically, for every communication round, each agent $i$ breaks its policy parameters $\theta^{(i)}$ into $P_{i}$ ($P_{i}=\min\{P,|\Omega_{i}|\}$, according to the default segmentation granularity $P$ and the number of neighbors $|\Omega_{i}|$ in current time) non-overlapping segments $\hat{\theta}^{(i)}_{1},\hat{\theta}^{(i)}_{2},\cdots,\hat{\theta}^{(i)}_{P_{i}}$ as
$\theta^{(i)}=(\hat{\theta}^{(i)}_{1},\hat{\theta}^{(i)}_{2},\cdots,\hat{\theta}^{(i)}_{P_{i}}){.}\vspace{-0.1em}$ (16)
@@
Significantly, the available segmentation strategies are diverse and include, but are not limited to, dividing the policy parameters by splitting DNN layers [33], modular approach [67], or segmenting according to the parameter size [66]. Without loss of generality, taking IoV as an application example, there exist rather diverse V2V communication protocols and technologies such as DSRC (based on IEEE 802.11p and its subsequent IEEE 802.11bd), NR-V2X (as specified in 3GPP Release 16 and 17), which support Multiple Input Multiple Output (MIMO) technology and advanced signal processing techniques, such as spatial multiplexing and beamforming. Therefore, the communication capability depends on multiple factors ranging from physical layer configurations, such as antenna array setup and channel conditions, to upper layer protocols, shaping the theoretical upper limit for segmentation granularity $P_{\max}$. Besides, shorter distances between agents (e.g., vehicles) and better channel conditions generally enhance the capacity and reliability of the system. Still, the segmentation also can be dynamic, with larger $P_{\max}$ in environments with higher agent density and smaller $P_{\max}$ in cases where fewer agents are within communication range.
@@
To clarify this process, we use the most intuitive uniform parameter partition. For each segment $p=1,\cdots,P_{i}$, agent $i$ randomly selects a target agent (without replacement) from its neighbors (i.e., $j_{p}\in\Omega_{i}$) to send segment request $(i,P_{i},j_{p},p)$, which indicates the agent $i$ who initiates the request, its total segment number $P_{i}$, as well as the requested segment $p$ from the target agent $j_{p}$. Upon receiving the request, the agent $j_{p}$ will break its own policy parameters $\theta^{(j_{p})}$ into $P_{i}$ segments and return the corresponding requested segment $\hat{\theta}^{(j_{p})}_{p}$ according to the identifier $p$. Then, agent $i$ could reconstruct a referential policy based on all of the fetched segments, that is,
$\tilde{\theta}^{(i)}=(\hat{\theta}^{(j_{1})}_{1},\hat{\theta}_{2}^{(j_{2})},\cdots,\hat{\theta}_{P_{i}}^{(j_{P_{i}})}){.}\vspace{-0.1cm}$ (17)
@@
In fact, this segmented transmission approach can be executed in parallel, thereby optimizing the utilization of available bandwidth. Instead of being confined to a single link, the traffic is distributed across $P_{i}$ links, enhancing the overall data transfer efficiency. Besides, in order to further accelerate the propagation and ensure the model quality, we can construct multiple model replicas in RSM-MASAC. That is, the process of segment request &\&& response can be repeated $P_{i}\times\kappa_{i}$ times, reconstructing $\kappa_{i}=\min\{\kappa,|\Omega_{i}|\}$ reconstructed referential policies in one communication round.
Algorithm 1 The RSM-MASAC Algorithm.
1:  Initialize network parameters $\theta^{(i)}$, $\omega_{1}^{(i)}$, $\omega_{2}^{(i)}$, $i=1,2,\cdots N$.
2:  Initialize target network parameters $\bar{\omega}_{1}^{(i)}\leftarrow\omega_{1}^{(i)}$, $\bar{\omega}_{2}^{(i)}\leftarrow\omega_{2}^{(i)}$.
@@
3:  Initialize learning rate $\eta_{\pi}$, $\eta_{Q}$, $\eta_{\alpha}$, temperature $\alpha$, communication interval $U$, segmentation granularity $P$, predefined replicas $\kappa$.
4:  Initialize iteration index $k\leftarrow 0$ and $counter\leftarrow 0$.
@@
9:        Select an action $a_{t}^{(i)}$ with respect to $s_{t}^{(i)}$ according to the current policy $\pi^{(i)}$ parameterized by $\theta^{(i)}$.
10:        Observe reward $r_{t}^{(i)}$ and next state $s_{t+1}^{(i)}$.
11:        Save the new transition in replay buffer: $\mathcal{D}^{(i)}\leftarrow\mathcal{D}^{(i)}\cup\langle s_{t}^{(i)},a_{t}^{(i)},r_{t}^{(i)},s_{t+1}^{(i)}\rangle$.
12:        Sample a mini-batch $\Phi^{(i)}\sim\mathcal{D}^{(i)}$.
13:        Update soft $Q$ function $\omega_{x}^{(i)}\leftarrow\omega_{x}^{(i)}-\eta_{Q}\nabla J_{Q}(\omega_{x}^{(i)})$ by (6), for $\forall x={1,2}$.
@@
15:           Update policy $\vspace{-0.5em}\theta_{k+1}^{(i)}\leftarrow\theta_{k}^{(i)}-\eta_{\pi}\nabla J_{\pi}(\theta_{k}^{(i)})$ by (7).
16:           Adjust temperature $\alpha^{(i)}\leftarrow\alpha^{(i)}-\eta_{\alpha}\nabla J(\alpha^{(i)})$ by (9).
17:           Update target networks $\bar{\omega}_{x}^{(i)}\leftarrow\varrho\omega_{x}^{(i)}+(1-\varrho)\bar{\omega}_{x}^{(i)}$, for $\forall x=1,2$.
18:           $k\leftarrow k+1$.
@@
20:        $counter\leftarrow counter+1$.
@@
23:           Update policy $\theta_{k}^{(i)}\leftarrow$ CommMix($\theta_{k}^{(i)}$, $i$, $\Omega_{i}$, $P$, $\kappa$) according to Algorithm 2.
@@
Algorithm 2 The CommMix Function in Algorithm 1.
1:  Input: $\theta$, $i$, $\Omega_{i}$, $P$, $\kappa$.
2:  $P_{i}=\min\{P,|\Omega_{i}|\}$ and $\kappa_{i}=\min\{\kappa,|\Omega_{i}|\}$.
3:  for each replica $1,2,\cdots,\kappa_{i}$ do
4:     Send $P_{i}$ pulling request $(i,P_{i},j_{p},p)$ to nearby collaborators in $\Omega_{i}$, and receive $\hat{\theta}^{(j_{p})}_{p}$ to reconstruct $\tilde{\theta}$ as (17).
5:     Select $M$ samples from the replay buffer $\mathcal{D}^{(i)}$.
6:     Estimate $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ according to (19).
7:     if $\mathbb{A}^{+}_{\pi}(\tilde{\pi})>0$ then
8:        Evaluate $F(\theta)$ according to (20).
9:        Get the upper bound of $\zeta$ according to Theorem 2.
10:        Make the mixture metric $\zeta$ less than the calculated upper bound, and update $\theta_{\text{mix}}^{(i)}$ by (10).
11:     end if
12:  end for
13:  Output: $\theta_{\text{mix}}^{(i)}$.
@@ V-A2 Policy Parameter Mixture with Theory-Established Performance Improvement
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\frac{1}{2}\zeta^{2}(\tilde{\theta}-\theta)^{\intercal}F(\theta)(\tilde{\theta}-\theta){,}\vspace{-2em}$
@@
$\vspace{-0.5em}F(\theta)=\mathop{\mathbb{E}}_{\begin{subarray}{c}s\sim d_{\pi}\\ a\sim\pi_{\theta}\end{subarray}}\left[\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)\!\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)^{\intercal}\right]{.}$
@@
$\qquad+\alpha[H(\tilde{\pi}_{\tilde{\theta}}(\cdot|s_{t}))-H({\pi}_{{\theta}}(\cdot|s_{t}))]\bigg{]}{,}$ (19)
@@
$F(\theta)\!\approx\!\mathop{\mathbb{E}}\limits_{s_{t}\sim\mathcal{D}}\!\bigg{[}\!\mathop{\mathbb{E}}\limits_{a_{t}\sim\pi_{\theta}}\!\bigg{[}\!\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)\!\left(\frac{\partial\log\pi_{\theta}(a|s)}{\partial\theta}\right)^{\intercal}\!\bigg{]}\bigg{]}{.}$ (20)
@@ V-B Communication Cost Analysis
Regarding the communication overhead for each segment request, RSM-MASAC incurs a maximization data transmission cost of $\upsilon_{p}=\upsilon/P$ through D2D communications. Following the same settings as in [67], we provide illustrative IoV examples to demonstrate the overall reduction in communication overhead. Specifically, the Collective Perception Message (CPM) defined in Collective Perception Service (CPS) of ETSI TR 103 562 [69] can encapsulate DNN parameters into the Perceived Object Containers (POCs) and propagate them among cooperating DRL agents with each message carrying a payload of $4,480$ bytes, serving as segment responses in our algorithm. And the DNN parameters $\theta\in\mathbb{R}^{d}$ ’s transmission size $\upsilon$ can be regarded as $32d$ bits, commonly assumed in [58, 67, 25]. Consequently, the total communication overhead for reconstructing maximal $\kappa$ referential policies per agent in each round $c(\upsilon,f)=\frac{N\times\upsilon\times\kappa}{8\times 1024^{3}}$ (GB), as well as $\frac{N\times\upsilon\times\kappa}{8\times 4480}$ message numbers, which is $(N-1)/\kappa$ times less than that in a fully connected communication setup [39]. Additionally, communication occurs at a periodic interval of $U$, allowing for further reduction in communication overhead by decreasing communication frequency. Moreover, by simultaneously requesting $P_{i}$ agents in parallel, RSM-MASAC benefits from the sufficient use of the bandwidth and enhances the capability to overcome possible channel degradation.
@@ V-C Complexity Analysis
For each referential policy, the calculation of policy advantage $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ is contingent with sample size $M$, policy parameter size $|d|$ and $Q$ network parameter size $|d_{Q}|$, with complexity $\mathcal{O}(M(|d|+|d_{Q}|))$. However, the mixture of policy parameters only proceeds if condition $\mathbb{A}^{+}_{\pi}(\tilde{\pi})>0$ is met during the communication-assisted mixing phase, with the computation complexity primarily depending on the calculation of FIM. Specifically, with sampling approximated method using first-order gradients in (20), the computation complexity of FIM is $\mathcal{O}(M|d|^{2})$ and the associated memory complexity is $\mathcal{O}(|d|^{2})$. However, the actual computation of the FIM is not restricted to sampling-based method alone, there has been extensive research on FIM approximation methods, including Diagonal Approximation [70, 71] and Low-Rank Approximation [72], among others, to further reduce to the complexity.
@@ VI Experimental Results and Discussions
In this section, we validate the effectiveness of our proposed algorithm for the speed control of Connected Automated Vehicles (CAVs) in IoV, highlighting its superiority compared to other methods.
@@ VI-A Experimental Settings
Parameters Definition Figure 8 Merge
@@
  • “Figure 8”: $14$ vehicles navigate a one-way lane shapes like a figure “8”, including $5$ emulated Human Driven Vehicles (HDV) controlled by Simulation of Urban MObility (SUMO) with Intelligent Driver Model (IDM) [75], and $9$ IRL-controlled CAVs maintaining dedicated links to update their parameters through the V2V channel. At the lane’s intersection, each CAV adjusts its acceleration to traverse efficiently, aiming to boost the traffic flow’s average speed.
  • “Merge”: A highway on-ramp merging scenario with vehicle flow at $2,300$ per hour, including a maximum of $2,200$ vehicles on the main road and $100$ vehicles on the ramp. Within each epoch, $13$ vehicles are randomly chosen to instantiate the DRL-based controllers as they sequentially enter, aiming to manage collision avoidance and congestion at merge points. The simulation settings closely align with those of the first scenario.
Both two scenarios are modified to assign the limited partial observation of the global environment as the state of each CAV, including the position and speed of its own, the vehicle ahead and behind. Only CAVs can execute the V2V end-to-end communication, and the connectivity of V2V links at time $t$ depends on the CAVs’ position and communication range, which are both extracted using the TraCI simulator in our experiments. A communication range of $90$ m is adopted in interactions “Figure 8” and $400$ m in highway “Merge”. This is according to the conclusion in [76] that communication range at intersections will be extremely reduced compared with the conventional scenarios [77], and relevant service requirements[78]. Unless otherwise stated, our experiments are based on the common MARL V2V lossless ideal communication premise [17, 27, 7, 28]. Meanwhile, each CAV’s action is a continuous variable representing speed acceleration or deceleration. It is sampled from the outputs of the policy network, which has three fully connected layers with $256$ hidden units each and ReLU activations. Actions are squashed using a tanh function to fall within $[-1,1]$ and then scaled by the acceleration bounds of the scenarios.
@@
Number of samples to evaluate $\mathbb{A}^{+}_{\pi}(\tilde{\pi})$ $M$ $50$
Learning rate of actor network $\eta_{\pi}$ $4\times 10^{-5}$
Learning rate of critic network $\eta_{Q}$ $3\times 10^{-4}$
@@
Segmentation granularity $P$ 4
Predefined replicas $\kappa$ 3
(a) Runtime on CPU
(b) Runtime on GPU
Figure 4: The runtime of estimating mixture metric under different sample and parameter sizes.
@@
For each referential policy, parameter mixing occurs only when $\mathbb{A}^{+}_{\pi}(\tilde{\pi})>0$. In this case, the computational complexity is predominantly governed by the calculation of $F(\theta)$ due to the outer product of policy gradients. Using an AMD EPYC 7F32 @$3.70$ GHz, $8$-core CPU and an NVIDIA RTX3090 GPU [^2] We note that there exist available automotive chips like the NVIDIA DRIVE Orin or Thor featuring CUDA Tensor Core GPU., as shown in Fig. 4, the computational time escalates with increasing sample size and parameter size, consistent with the discussions in Section V-C. However, with the aid of GPU hardware acceleration, such as superior parallel processing capabilities, higher memory bandwidth, and optimized deep learning libraries, the runtime experiences a significant reduction.
In order to reduce the occurrence of collisions and promote the traffic flow to the maximum desired speed, we take the normalized average speed of all vehicles at each timestep as the individual reward in each scenario [^3] Notably, we assume complete knowledge of individual vehicle speeds at each vehicle here. Beyond the scope of this paper, some value-decomposition methods like [79] can be further leveraged to derive a decomposed reward, so as to loosen such a strict requirement., which is assigned to each training agent after its action is performed. In addition, the current epoch will be terminated once a collision occurs or the max length of step $T$ in an epoch is reached.
The results with variance are averaged over $5$ independent simulations, with tests conducted every $10$ epochs during training. Moreover, as for the baseline, we take the vehicles controlled by the Flow IDM [75], which belongs to a typical car-following model incorporating extensive prior knowledge and indicates the pinnacle of performance achievable by the best centralized federated MARL algorithms [14]. Furthermore, the principal hyper-parameters used in simulations are listed in Table IV.